OpenStudy (anonymous):

Five waves pass a dock in 2.0 s. If the waves are traveling at 10.0 m/s, then what is the wavelength of the waves?

OpenStudy (anonymous):

v = d/t d = v*t d = 20m/s 5 waves Answer = 20meters/5 4 ....?

OpenStudy (roadjester):

\(\Huge \lambda f=c\) Where lambda is the wavelength, f is the frequency, and c is the speed of light.

OpenStudy (roadjester):

hmm, based on the information given, I can't assume that it's an electromagnetic wave. Therefore instead of using "c", use "v", the velocity.

OpenStudy (roadjester):

\(\Huge \lambda f=v\)

OpenStudy (anonymous):

wavelength times frequency = velocity?

OpenStudy (roadjester):

yep

OpenStudy (anonymous):

wavelength times 2.0s = 10 m/s?

OpenStudy (roadjester):

(wavelength)(#of waves per time)=velocity

OpenStudy (anonymous):

5/2 would be the waves per time?

OpenStudy (roadjester):

yup, so for this simple case, your original calculation matches. For more complex waves, what you didn't would be less likely to work out

OpenStudy (anonymous):

so it would be wavelength times 2.5 = 10.0 m/s?

OpenStudy (roadjester):

yup

OpenStudy (anonymous):

so the wavelength would be 4 m?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

oh! just one last question sorry I completely forgot! " In a thunderstorm at 20.0°C, Amanda sees a bolt of lightning and hears the thunderclap 3.00 s later. How far from Amanda did the lightning strike?

OpenStudy (roadjester):

That's two questions. You passed ur own quota =) Bye bye.

OpenStudy (anonymous):

lol ok night!